Is the function 𝑓:𝑹 →R defined as f:(x) = (x — 7) (x^3 + 11)is an odd function.

1
Expert's answer
2021-11-01T13:28:37-0400

"f(x) = (x - 7) (x^3 + 11)
\\f(-x) = (-x-7)[(-x)^3+11)]
\\=-(x+7)(-x^3+11)
\\=-(x+7)(11-x^3)\neq-f(x)"

For "f(x)" to be an odd function: "f(-x)=-f(x)"

But, there is no relation between "f(x)" and "f(-x)" here in this problem.

So, "f(x)" is not an odd function.